# Improving Molecule Generation and Drug Discovery with a Knowledge-enhanced Generative Model

### v1

@@ Abstract
Recent advancements in generative models have established state-of-the-art benchmarks in the generation of molecules and novel drug candidates. Despite these successes, a significant gap persists between generative models and the utilization of extensive biomedical knowledge, often systematized within knowledge graphs, whose potential to inform and enhance generative processes has not been realized. In this paper, we present a novel approach that bridges this divide by developing a framework for knowledge-enhanced generative models called K-DReAM. We develop a scalable methodology to extend the functionality of knowledge graphs while preserving semantic integrity, and incorporate this contextual information into a generative framework to guide a diffusion-based model. The integration of knowledge graph embeddings with our generative model furnishes a robust mechanism for producing novel drug candidates possessing specific characteristics while ensuring validity and synthesizability. K-DReAM outperforms state-of-the-art generative models on both unconditional and targeted generation tasks.
Drug discovery, Knowledge graphs, Generative models
Figure 1: K-DReAM a. A biomedical knowledge graph showing the links between drugs, target proteins, diseases, genes, biological pathways, etc. The edges represent different types of relationships like targeting, synergistic or unwanted interactions, or side effects. b. Given a relation and a subject, we can complete the triple and obtain Knowledge Graph Embeddings (KGEs) (Sec 4). c. These embeddings are used to guide a generative process (Sec 5). d. The end result of the generative process is a novel drug candidate.
1 Introduction
Drug discovery is an expensive endeavor, with costs often surging beyond the billion-dollar mark, primarily due to the extensive stages involved in the development pipeline, ranging from target identification to clinical evaluations (Morgan et al., 2011). Nevertheless, the advent of computational techniques and machine learning has ushered in a paradigm shift in pharmaceutical research, streamlining drug development, mitigating expenses, and enhancing the discovery of bioactive compounds (Rudrapal & Chetia, 2020). Predominantly, the emergence of sophisticated machine learning applications has substantially bolstered predictive capabilities, accelerating the pace of in-silico drug design (Rifaioglu et al., 2018). However, this rapid progression has not been devoid of challenges. Present-day generative models in machine learning showcase impressive results but often miss out on tapping into the extensive biological knowledge available in the domain (Tian et al., 2023). These models frequently grapple with issues like overfitting, lack of generalizability to new data (Ma et al., 2015), and difficulties in handling novel drug or target interactions (Nguyen et al., 2022). Although there is a vast amount of biomedical knowledge, there is a missing link between these collective datasets and the efforts to build generative models for drug discovery.
@@ 1 Introduction
Molecular generation is a critical aspect of drug discovery, material science, and chemical exploration. Generative models have demonstrated the ability to generate novel structures across a variety of representations, from 1-D SMILES strings to 2-D and 3-D molecular graphs. However, while these models have shown great results in producing drug-like molecules, they are typically evaluated based on their unconditioned generation capabilities. Generative models have also achieved significant improvements in targeted generation with modalities like text, images (Black et al., 2023), sound, and videos. The development of a targeted generative model for graphs is of particular interest in the problem of drug discovery.
Biomedical knowledge can be organized and represented by a powerful tool called a Knowledge Graph (KG). A good way of guiding the generative process is by using the KG as a structural and semantic scaffold, which can provide context-relevant constraints and objectives for the generative models. This integration of KGs with generative models can facilitate the targeted generation of molecules by driving the model to consider biological relevance and plausibility in the generation process. By doing so, the models can leverage both the relational knowledge embedded within the KG, such as the interactions between proteins, genes, drugs, and diseases, and the generative capabilities to propose novel compounds that are more likely to exhibit desired therapeutic effects and pharmacological profiles.
@@
We propose an end-to-end framework 1 for Knowledge-enhanced Drug discovery with a Reinforcement learning-Augmented Model (K-DReAM), with the following key contributions:
@@
    A diffusion-based generative model for graphs steered with Discrete Diffusion Policy Optimization (DDPO) for improved molecule generation. The generative model, initially trained on extensive chemical structure datasets, is fine-tuned with the DDPO approach that rewards the model for producing chemically valid, pharmacologically active structures that have the ability to satisfy user-defined properties. (Section 3)
@@
    A Knowledge Graph Embedding (KGE) model that generates contextual embeddings from large-scale biomedical databases that are used for targeted extrapolation and synthesis of drug candidates. By reformulating models that generate KGEs from energy-based models to ones that minimize a Maximum Likelihood Estimation (MLE) objective, we demonstrate the ability to generate more meaningful embeddings that also incorporate domain constraints, which ensure that predictions follow user-specified logical formulae. (Section 4)
@@
    The development of a training workflow that is used by the model to learn conditional generation, with the help of a property inference network that integrates the knowledge represented by the KGEs into the training of the diffusion model to create a unified process for knowledge-enhanced novel drug discovery (Section 5)
@@
2 Related Work
Molecular Generation. The use of SMILES strings (Weininger, 1988), a one-dimensional notation that encodes molecular structures into strings of ASCII characters, was popularized for generative modeling with the advent of recurrent neural networks (RNNs) and variational autoencoders (VAEs). (Gómez-Bombarelli et al., 2018) demonstrated that VAEs could be trained on SMILES strings to interpolate in latent space and generate molecules with desired chemical properties. Subsequent work by (Neil et al., 2018) employed RNNs for a similar purpose and introduced reinforcement learning to bias the generation process towards certain properties. While the SMILES-based approach remains popular due to its simplicity and the wealth of available chemical data, it is not without limitations, as SMILES strings do not explicitly capture the geometrical and topological features of molecules, and small changes in SMILES strings can lead to vastly different molecular structures (Lavecchia, 2019).
To address the limitations of SMILES, researchers have increasingly turned to graph-based representations, where molecules are represented as graphs with atoms as nodes and bonds as edges. This representation aligns with the intrinsic structure of molecules. Generative models have been devised using GANs (Martinkus et al., 2022), diffusion (Jo et al., 2022a; Vignac et al., 2023), autoregressive methods (Kong et al., 2023), and normalizing flows (Shi et al., 2019; Zang & Wang, 2020).
Knowledge-enhanced Models. Knowledge graphs have gained traction in drug discovery (James & Hennig, 2023) by structuring diverse biological data into interconnected frameworks that facilitate the identification of new therapeutic targets and drug repurposing opportunities (Santos et al., 2022; Zheng et al., 2020; Chandak et al., 2023). These graphs represent entities such as molecules, proteins, and diseases as nodes, with edges capturing their intricate relationships. (Yu et al., 2021) showcase the use of entity embeddings from knowledge graphs for predicting drug-disease associations, while (Zitnik et al., 2018) employed graph neural networks on similar structures to uncover drug-target interactions, surpassing traditional predictive models. By incorporating relational information from knowledge graphs, molecular generative models can impose biologically relevant constraints during the generation process (Bilodeau et al., 2022). Knowledge graphs also aid in systematically identifying adverse drug reactions (Nováček & Mohamed, 2020) and repurposing existing drugs by analyzing the network’s topology to reveal hidden biological pathways (Pan et al., 2022).
Diffusion Models. Diffusion models have emerged as a powerful class of generative models with applications in drug discovery, offering an alternative to traditional variational autoencoders and generative adversarial networks. They work by gradually introducing noise to a data distribution and learning to reverse this process, thereby generating new data points. In the context of drug discovery, diffusion models can generate novel molecular structures by learning the distribution of drug-like molecules, with Ho et al. (2020) demonstrating their potential through the generation of high-fidelity images. Specifically, these models can potentially be adapted to generate 2D molecular structures that could lead to novel compounds with desired properties.
3 Molecular Generative Model
@@
Our generative model is built on the foundations of the GDSS (Jo et al., 2022a) and MOOD (Lee et al., 2022) diffusion models. Graph Diffusion via the System of SDEs (GDSS) defines the forward diffusion $q$ of a graph $\textbf{G}_{t}=(\textbf{X}_{t},\textbf{A}_{t})$ with a Stochastic Differential Equation (SDE):
@@
(Black et al., 2023) introduce a fine-tuning technique known as Denoising Diffusion Policy Optimization (DDPO) for image-based models. This method excels in handling images by incorporating human feedback to ensure metrics like aesthetic qualities—a factor typically difficult to quantify in computational models. Despite its success in the domain of images, extending these techniques to the realm of molecule optimization presents significant challenges. The intricate dependencies between nodes and edges in molecular structures demand careful consideration, as these relationships are critical to determining the validity and inherent properties of the molecules.
We introduce a reward function $r$ that simultaneously prioritizes multiple molecular properties: (i) The drug-likeness $Q$ of a molecule, quantified by its QED score (Bickerton et al., 2012), (ii) the synthesizability $S$ calculated using the SAScore (Ertl & Schuffenhauer, 2009) which is a rule-based determination of synthetic accessibility, and (iii) a provision for defining a property $C$ in the chemical space like novelty/molecular similarity, docking scores, structure validity, etc.
@@
We elaborate on the training process of the model for conditional generation in Section 5.
4 Knowledge Graph Embeddings
A knowledge graph $\mathcal{G}$, involving a set of entities $\mathcal{E}$ and relations $\mathcal{R}$, is a directed multigraph composed of triples $(s,r,o)\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ that represent the relation $r$ between subject $s$ and object $o$ nodes. In biomedical knowledge graphs, $\mathcal{E}$ contains entities like drugs, diseases, genes, phenotypes, and biological pathways while $\mathcal{R}$ contains relations like interactions, drug targets and side effects. Some examples of public domain knowledge graphs in this field like Hetionet (Himmelstein et al., 2017), CKG (Santos et al., 2022), BioKG (Walsh et al., 2020), PharmKG (Zheng et al., 2020), and PrimeKG (Chandak et al., 2023), which range from $\sim$10K to 10M nodes and up to $\sim$200M relations.
@@ 4 Knowledge Graph Embeddings
KGE models involve various translational distance-based metrics like TransE (Lin et al., 2015), tensor factorization methods such as RESCAL (Nickel et al., 2011), and neural network-based models like ConvE (Dettmers et al., 2018). These models differ in how they interpret the relationships and interactions between entities, and each comes with relative strengths and weaknesses with regard to specific types of relational data and structures.
Traditionally, score-based KGE models have been interpreted as energy-based models, where the score is seen as a measure of the negative energy of an $(s,r,o)$ triplet. Recasting these models into a probabilistic interpretation would enable exact training by Maximum Likelihood Estimation (MLE), as well as the ability to encode domain constraints into the learning process. In order to transform these negative energies into probabilities over the $\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ space, we need to calculate the partition function, which is infeasible for large-scale biomedical KGs. (Loconte et al., 2023) demonstrate that KGE models can be represented by structured computational graphs, called circuits, which are expressive probabilistic functions over the triplet space and can be efficiently trained with MLE. This enables us to use KGE models as efficient generative models of new $(s,r,o)$ triples, consistent with the statistics of existing the KG while guaranteeing the satisfaction of constraints that would be crucial to pharmacological applications.
@@
Circuit-based score functions bring down the complexity of evaluating $Z$ from $\mathcal{O}(|\mathcal{E}|^{2}\cdot|\mathcal{R}|\cdot\text{cost}(\phi))$ to $\mathcal{O}(|\mathcal{E}+\mathcal{R}|\cdot\text{cost}(\phi))$ (Loconte et al., 2023).
The generation of new $(s,r,o)$ triplets is ensured to be semantically coherent within the rules of the KG by introducing domain constraints (Ahmed et al., 2022) that modify the score functions with indicator functions for valid triples. Given a relation $r\in\mathcal{R}$ the subsets $S_{r},O_{r}\subset\mathcal{E}$ represent the sets of entities between which $r$ is a valid relation. This domain $K_{r}=S_{r}\wedge\{r\}\wedge O_{r}$ can be extended across all $r$ to form $K=\vee_{r}K_{r}$. We restrict the score function to only valid triples $(s,r,o)\in K$ by introducing an indicator function $c_{K}$ in the score calculation $\phi(s,r,o)=p(s,r,o)\cdot c_{K}(s,r,o)$. For example, in our experiments, we are particularly interested in drug-protein interaction. For this type of relation $r=\texttt{targets}$, we restrict $s\in\texttt{drugs}$ and $o\in\texttt{proteins}$. Once again, the circuit-based score functions help us train these models with a complexity of $\mathcal{O}(|\mathcal{E}+\mathcal{R}|\cdot\text{cost}(\phi)\cdot\text{cost}(c_{K}))$
In order to generate KGEs, K-DReAM uses the RotatE (Sun et al., 2019) algorithm with the above modifications to the training process in order to restrict the model to generate coherent triples. The embeddings generated from the knowledge graph are used to guide the generative process as described below.
5 Implementation of K-DReAM
@@
5.1 Property Inference Network $P_{\phi}(\textbf{G})$
To guide the conditional generation process, we create a neural network to estimate knowledge-based embeddings c from a noised version of an input molecular graph $\textbf{G}_{T}$. $P_{\phi}(\textbf{G}_{T})\approx\textbf{c}$ is used to implement a modified version of the classifier guidance algorithm by (Sohl-Dickstein et al., 2015). While previous work by (Lee et al., 2022) and (Vignac et al., 2023) uses a similar algorithm to guide conditional generative processes, ours is a novel approach that utilizes a combination of graph attention and convolutional layers to estimate knowledge-based embeddings, effectively creating a map between chemical space and KGE space.

### v2

@@ Abstract
Recent advancements in generative models have established state-of-the-art benchmarks in the generation of molecules and novel drug candidates. Despite these successes, a significant gap persists between generative models and the utilization of extensive biomedical knowledge, often systematized within knowledge graphs, whose potential to inform and enhance generative processes has not been realized. In this paper, we present a novel approach that bridges this divide by developing a framework for knowledge-enhanced generative models called KARL. We develop a scalable methodology to extend the functionality of knowledge graphs while preserving semantic integrity, and incorporate this contextual information into a generative framework to guide a diffusion-based model. The integration of knowledge graph embeddings with our generative model furnishes a robust mechanism for producing novel drug candidates possessing specific characteristics while ensuring validity and synthesizability. KARL outperforms state-of-the-art generative models on both unconditional and targeted generation tasks.
Figure 1: KARL Model A Biomedical knowledge graph showing the links between drugs, target proteins, diseases, genes, biological pathways, etc. The edges represent different types of relationships like targeting, synergistic or unwanted interactions, or side effects. B We obtain Knowledge Graph Embeddings (KGEs) (Sec IV) for the target drug candidate. C These embeddings are used to guide a generative process (Sec V). D The end result of the generative process is a novel drug candidate.
I Introduction
Drug discovery is an expensive endeavor, with costs often surging beyond the billion-dollar mark, primarily due to the extensive stages involved in the development pipeline, ranging from target identification to clinical evaluations [1]. Nevertheless, the advent of computational techniques and machine learning has ushered in a paradigm shift in pharmaceutical research, streamlining drug development, mitigating expenses, and enhancing the discovery of bioactive compounds [2]. Predominantly, the emergence of sophisticated machine learning applications has substantially bolstered predictive capabilities, accelerating the pace of in-silico drug design [3]. However, this rapid progression has not been devoid of challenges. Present-day generative models in machine learning showcase impressive results but often miss out on tapping into the extensive biological knowledge available in the domain [4]. These models frequently grapple with issues like overfitting, lack of generalizability to new data [5], and difficulties in handling novel drug or target interactions [6]. Although there is a vast amount of biomedical knowledge, there is a missing link between these collective datasets and the efforts to build generative models for drug discovery.
@@ I Introduction
Molecular generation is a critical aspect of drug discovery, material science, and chemical exploration. Generative models have demonstrated the ability to generate novel structures across a variety of representations, from 1-D SMILES strings to 2-D and 3-D molecular graphs. However, while these models have shown great results in producing drug-like molecules, they are typically evaluated based on their unconditioned generation capabilities. Generative models have also achieved significant improvements in targeted generation with modalities like text, images [7], sound, and videos. The development of a targeted generative model for graphs is of particular interest in the problem of drug discovery.
Biomedical knowledge can be organized and represented by a powerful tool called a Knowledge Graph. A good way of guiding the generative process is by using the KG as a structural and semantic scaffold, which can provide context-relevant constraints and objectives for the generative models. This integration of KGs with generative models can facilitate the targeted generation of molecules by driving the model to consider biological relevance and plausibility in the generation process. By doing so, the models can leverage both the relational knowledge embedded within the KG, such as the interactions between proteins, genes, drugs, and diseases, and the generative capabilities to propose novel compounds that are more likely to exhibit desired therapeutic effects and pharmacological profiles.
@@
We propose an novel model capable of end-to-end generation (Figure 1) called the Knowledge-Augmented Reinforcement Learning-optimized (KARL) Model. This paper presents the following key innovations:
@@
    A diffusion-based generative model for graphs steered with Discrete Diffusion Policy Optimization (DDPO) for improved molecule generation. The generative model, initially trained on extensive chemical structure datasets, is fine-tuned with the DDPO approach that rewards the model for producing chemically valid, pharmacologically active structures that have the ability to satisfy user-defined properties. (Section III)
@@
    A Knowledge Graph Embedding (KGE) model that generates contextual embeddings from large-scale biomedical databases that are used for targeted extrapolation and synthesis of drug candidates. By reformulating models that generate KGEs from energy-based models to ones that minimize a Maximum Likelihood Estimation (MLE) objective, we demonstrate the ability to generate more meaningful embeddings that also incorporate domain constraints, which ensure that predictions follow user-specified logical formulae. (Section IV)
@@
    The development of a training workflow that is used by the model to learn conditional generation, with the help of a property inference network that integrates the knowledge represented by the KGEs into the training of the diffusion model to create a unified process for knowledge-enhanced novel drug discovery (Section V)
@@
II Related Work
Molecular Generation.
Molecular representations can be classified into three broad categories: (i) string-based (1-D), (ii) graph-based (2-D), and coordinate-based (3-D) representations. The simplified molecular-input line-entry system (SMILES) [8] remains one of the most widespread implementations of the 1-D representation. SMILES has been used with a variety of models such as Variational Auto-encoders (VAEs) [9] [10], Generative Adversarial Networks (GANs) like [11], and Recurrent Neural Networks (RNNs) [12]. The simplified nature of SMILES strings does not explicitly capture geometrical and topological features. Additionally, small changes in the strings can lead to vastly different or invalid molecular structures [13], which is of considerable concern while using diffusion-based models. In order to mitigate these problems, alternative models like [14] [15] using the SELF-referencIng Embedded Strings (SELFIES) [16] have been developed which ensure validity and complete coverage of chemical space.
Two-dimensional representations present the opportunity of directly working with the structure of compounds. Although they do not take into account the 3-D structure of the molecule, 2-D graph-based representations have achieved impressive results using models like GANs [17], normalizing flows [18] [19], and diffusion-based models [20] [21] [22] [23].
Knowledge-enhanced Models.
Knowledge graphs have gained traction in drug discovery [24] by structuring diverse biological data into interconnected frameworks that facilitate the identification of new therapeutic targets and drug repurposing opportunities [25], [26], [27]. These graphs represent entities such as molecules, proteins, and diseases as nodes, with edges capturing the connections and associations between them by encoding predicates like ‘causes,’ ‘indicates,’ ‘interacts,’ ‘targets,’ etc. The paper by [28] showcases the use of entity embeddings from knowledge graphs for predicting drug-disease associations, while [29] employed graph neural networks on similar structures to uncover drug-target interactions, surpassing traditional predictive models. By incorporating relational information from knowledge graphs, molecular generative models can impose biologically relevant constraints during the generation process [30]. Knowledge graphs also aid in systematically identifying adverse drug reactions [31] and repurposing existing drugs by analyzing the network’s topology to reveal hidden biological pathways [32].
III Molecular Generative Model
@@
Our generative model is built on the foundations of the GDSS [20] and MOOD [33] diffusion models. Graph Diffusion via the System of SDEs (GDSS) defines the forward diffusion $q$ of a graph $\textbf{G}_{t}=(\textbf{X}_{t},\textbf{A}_{t})$ with a Stochastic Differential Equation (SDE):
@@
Black et al. [7] introduce a fine-tuning technique known as Denoising Diffusion Policy Optimization (DDPO) for image-based models. This method excels in handling images by incorporating human feedback to ensure metrics like aesthetic qualities—a factor typically difficult to quantify in computational models. Despite its success in the domain of images, extending these techniques to the realm of molecule optimization presents significant challenges. The intricate dependencies between nodes and edges in molecular structures demand careful consideration, as these relationships are critical to determining the validity and inherent properties of the molecules.
We introduce a reward function $r$ that simultaneously prioritizes multiple molecular properties: (i) The drug-likeness $Q$ of a molecule, quantified by its QED score [34], (ii) the synthesizability $S$ calculated using the SAScore [35] which is a rule-based determination of synthetic accessibility, and (iii) a provision for defining a property $C$ in the chemical space like novelty/molecular similarity, docking scores, structure validity, etc.
@@
We elaborate on the training process of the model for conditional generation in Section V.
IV Knowledge Graph Embeddings
A knowledge graph $\mathcal{G}$, involving a set of entities $\mathcal{E}$ and relations $\mathcal{R}$, is a directed multigraph composed of triples $(s,r,o)\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ that represent the relation $r$ between subject $s$ and object $o$ nodes. In biomedical knowledge graphs, $\mathcal{E}$ contains entities like drugs, diseases, genes, phenotypes, and biological pathways while $\mathcal{R}$ contains relations like interactions, drug targets and side effects. Some examples of public domain knowledge graphs in this field like Hetionet [36], CKG [25], BioKG [37], PharmKG [26], and PrimeKG [27], which range from $\sim$10K to 10M nodes and up to $\sim$200M relations.
@@ IV Knowledge Graph Embeddings
KGE models involve various translational distance-based metrics like TransE [38], tensor factorization methods such as RESCAL [39], and neural network-based models like ConvE [40]. These models differ in how they interpret the relationships and interactions between entities, and each comes with relative strengths and weaknesses with regard to specific types of relational data and structures.
Traditionally, score-based KGE models have been interpreted as energy-based models, where the score is seen as a measure of the negative energy of an $(s,r,o)$ triplet. Recasting these models into a probabilistic interpretation would enable exact training by Maximum Likelihood Estimation (MLE), as well as the ability to encode domain constraints into the learning process. In order to transform these negative energies into probabilities over the $\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ space, we need to calculate the partition function, which is infeasible for large-scale biomedical KGs. [41] demonstrate that KGE models can be represented by structured computational graphs, called circuits, which are expressive probabilistic functions over the triplet space and can be efficiently trained with MLE. This enables us to use KGE models as efficient generative models of new $(s,r,o)$ triples, consistent with the statistics of existing the KG while guaranteeing the satisfaction of constraints that would be crucial to pharmacological applications.
@@
Circuit-based score functions bring down the complexity of evaluating $Z$ from $\mathcal{O}(|\mathcal{E}|^{2}\cdot|\mathcal{R}|\cdot\text{cost}(\phi))$ to $\mathcal{O}(|\mathcal{E}+\mathcal{R}|\cdot\text{cost}(\phi))$ [41].
The generation of new $(s,r,o)$ triplets is ensured to be semantically coherent within the rules of the KG by introducing domain constraints [42] that modify the score functions with indicator functions for valid triples. Given a relation $r\in\mathcal{R}$ the subsets $S_{r},O_{r}\subset\mathcal{E}$ represent the sets of entities between which $r$ is a valid relation. This domain $K_{r}=S_{r}\wedge\{r\}\wedge O_{r}$ can be extended across all $r$ to form $K=\vee_{r}K_{r}$. We restrict the score function to only valid triples $(s,r,o)\in K$ by introducing an indicator function $c_{K}$ in the score calculation $\phi(s,r,o)=p(s,r,o)\cdot c_{K}(s,r,o)$. For example, in our experiments, we are particularly interested in drug-protein interaction. For this type of relation $r=\texttt{targets}$, we restrict $s\in\texttt{drugs}$ and $o\in\texttt{proteins}$. Once again, the circuit-based score functions help us train these models with a complexity of $\mathcal{O}(|\mathcal{E}+\mathcal{R}|\cdot\text{cost}(\phi)\cdot\text{cost}(c_{K}))$.
In order to generate KGEs (Fig 1 C), KARL uses the RotatE [43] algorithm with the above modifications to the training process in order to restrict the model to generate coherent triples. The embeddings generated from the knowledge graph are used to guide the generative process as described below.
V Implementation of KARL
@@
V-A Property Inference Network $P_{\phi}(\textbf{G})$
To guide the conditional generation process, we create a neural network to estimate knowledge-based embeddings c from a noised version of an input molecular graph $\textbf{G}_{T}$. $P_{\phi}(\textbf{G}_{T})\approx\textbf{c}$ is used to implement a modified version of the classifier guidance algorithm by [44]. While previous work by [33] and [21] uses a similar algorithm to guide conditional generative processes, ours is a novel approach that utilizes a combination of graph attention and convolutional layers to estimate knowledge-based embeddings, effectively creating a map between chemical space and KGE space.